OpenStudy (anonymous):

please help me understand!!! :) new lesson that I CANNOT seem to grasp... :( Simplify √-121

OpenStudy (anonymous):

My answer: -11 (upside down exclamation point)

hartnn (hartnn):

\(\huge \sqrt{-1}=i\)

OpenStudy (anonymous):



hartnn (hartnn):

\(\sqrt{-121}=\sqrt{-1}\sqrt{121}=?\)

OpenStudy (anonymous):

Not quite sure, dear

hartnn (hartnn):

 now ?